# Telegrapher’s Generative Model via Kac Flows

### v3

@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
Denote the conditional velocity field corresponding to $u_{t}(\cdot|x_{0})$ by $v_{t}(\cdot|x_{0})$. Then, the vector field
@@ 7.2 Image Generation
In general, the ’standard’ Kac process $X_{t}=X_{0}+K_{t}$ and $K_{t}$ themselves are not equal at the final time $T$. It is common practice in generative modeling to augment the process so that the data no longer contributes to the final distribution (compare with variance-preserving diffusion processes, e.g. the Ornstein–Uhlenbeck process). Additionally, we introduce a time-scheduling function for the noising process. For simplicity, we fix $T=1$ in the rest of this section.
The Mean-Reverting Kac Process.
We define the mean-reverting Kac process by
@@
where $v_{K}(t,x\mid 0)$ is the conditional Kac velocity field from (86). By the same arguments as in Corollary 14, one recovers the corresponding vector field for the mean-reverting Kac process.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". R. Duong, P. Friz and G. Steidl acknowledge funding of the project EF-LI-Opt-2 within the DFG Excellence Cluster MATH+. The work of P. Friz is also supported by DFG CRC/TRR 388, project A02. We thank Gregor Kornhardt, Matthias Liero and Nicolaj Rux for fruitful discussions.
@@ A.1 Simulating the Kac Process
This procedure can be applied independently for multiple values of $t$, allowing efficient simulation of the Kac process. See Algorithm 1.
@@ A.2 Implementation Details
Algorithm 2 Sampling the 1D Kac density via inverse-CDF
@@
7:  Compute $x\;\approx\;F_{\rm cont}^{-1}(V;\,t)$, where $F_{\rm cont}$ is the continuous-part CDF in (32), numerically approximated via a precomputed inverse-CDF table.
@@
where $x^{(n)}$ are drawn from the learned model. We set $N=5000$. For the velocity field we used a fully-connected MLP with three hidden layers of width $w=256$ and ReLU activations. Specifically for the diffusion model we required time truncation, namely we train the flow on the interval $\varepsilon,\,T]$ with $\varepsilon=10^{-15}$. We used the torchdiffeq [8] package to simulate the ODEs with "dopri5" adaptive step size solver.

### v4

@@ 6 Velocity Field of the Multi-Dimensional Kac Flow
Denote the conditional velocity field corresponding to $u_{t}(\cdot|x_{0})$ by $v_{t}(\cdot|x_{0})$. By independence of $\mathbf{X}_{0}$ and $\mathbf{K}(t)$, it is given via (91), and finally, the vector field
@@ 7.2 Image Generation
In general, the ’standard’ Kac process $X_{t}=X_{0}+K_{t}$ and $K_{t}$ themselves are not equal at the final time $T$. It is common practice in generative modeling to augment the process so that the data no longer contributes to the final distribution (compare with variance‐preserving diffusion processes, e.g. the Ornstein–Uhlenbeck process). Additionally, we introduce a time‐scheduling function for the noising process. Let $X_{0}$ and $K(t)$ be independent, and for simplicity, we fix $T=1$ for the rest of this section.
The Mean‐Reverting Kac Process.
We define the mean‐reverting Kac process by
@@
where $v_{K}(t,x\mid 0)$ is the conditional Kac velocity field from (86). Note that in the last equality, the independence of $X_{0}$ and $K(t)$ is used to drop the condition $X_{0}=x_{0}$. Now, by the same arguments as in Corollary 14, one recovers the corresponding vector field for the mean‐reverting Kac process.
@@ Acknowledgement.
J. Chemseddine acknowledges funding of the DFG project STE 571/17-2 within the SPP 2298 "Theoretical Foundations of Deep Learning". R. Duong, P. Friz and G. Steidl acknowledge funding of the project EF-LI-Opt-2 within the DFG Excellence Cluster MATH+. The work of P. Friz is also supported by DFG CRC/TRR 388, project A02. Our thanks go to Gregor Kornhardt, Matthias Liero, Nicolaj Rux, and Weiqiao Han for fruitful discussions. We are especially grateful to Weiqiao Han for correcting an error in our implementation.
@@ A.1 Simulating the Kac Process
This procedure can be applied independently for multiple values of $t$, allowing efficient simulation of the Kac process, see Algorithm 1.
@@ A.2 Implementation Details
Algorithm 2 Sampling the 1D Kac density via inverse‐CDF
@@
7:  Compute $x\;\approx\;F_{\rm cont}^{-1}(V;\,t)$, where $F_{\rm cont}$ is the continuous‐part CDF in (32), numerically approximated via a precomputed inverse‐CDF table.
@@
where $x^{(n)}$ are drawn from the learned model. We set $N=5000$. For the velocity field we used a fully‐connected MLP with three hidden layers of width $w=256$ and ReLU activations. Specifically for the diffusion model we required time truncation, namely we train the flow on the interval $\varepsilon,\,T]$ with $\varepsilon=10^{-15}$. We used the torchdiffeq [8] package to simulate the ODEs with "dopri5" adaptive step size solver.